VIBRATIONALLY MEDIATED PHOTODISSOCIATION OF HCN

\underline{C. R. Bucher}, B. M. Harvey AND K. K. Lehmann, Department of Chemistry, Princeton University, Princeton N. J. 08544.

Recently, the dissociation mechanism of HCN by vibrationally mediated photodissociation has been examined. Exciting the C-H stretch with four quanta and photodissociating the molecule with 220 nm light generates CN in the first excited electronic Ã ^{2}Π state rather than the ground X̃ ^{2}Σ^{+} state. The CN fragment is detected by probing the B̃ ^{2}Σ^{+} ← Ã ^{2}Π (1-0) band and monitoring the laser induced fluorescence (LIF) via the B̃ ^{2}Σ^{+} → X̃ ^{2}Σ^{+} (1-1) transition. Progress on the analysis of the LIF spectrum and the nascent product distribution will be presented.